A passanger train can travel 325 mi in the same time a freight train takes to travek 200 mi. If the speed of the passanger train is 25 mi.h faster then the speed of the freight train, find the speed of each.
Answer by checkley75(3666)

TIME=DISTANCE/RATE. SEEING THAT THE TIMES ARE EQUAL THEN WE HAVE THE FOLLOWING EQUATION:
325/(R+25)=200/R NOW CROSS MULTIPLY:
325R=200(R+25)
325R=200R+5000
325R-200R=5000
125R=5000
R=5000/125
X=40 MPH FOR THE FREIGHT TRAIN.
PROOF
325(40+25)=200/40
325/65=200/40
5=5